Us four unit multipliers to convert 20 square feet to square centimeters
Answer by Alan3354(69443)

1 ft = 12 inches = 12*2.54 cm = 30.48 cm
20 sq ft = 20*30.48^2 sq cm
= 18580.608 sq cm